# ActSonic: Everyday Activity Recognition on Smart Glasses using Active Acoustic Sensing

### v1

@@ 8.6. Potential Real-world Application
The promising performance of ActSonic recognizing 27 activities in the wild using low-power and minimally-obtrusive glasses will significantly lower the barriers to logging everyday activities. It would further create opportunities for many downstream applications that are based on tracking one or multiple types of activities. Here we list a few sample applications :
@@ 8.6.2. High-Resolution Behavior Data in the Wild for Health
ActSonic can track 27 everyday activities, many of which are related to health behaviors. Automatically logging these health-related activities can potentially provide researchers and clinical physicians an opportunity to better understand a user’s activities in the wild for health purposes. For instance, the eating and drinking behavior can potentially be used to analyze the user’s eating and drinking routines related to eating disorders and body hydration levels. Brushing teeth, flossing, and rinsing the mouth can be recognized with a very low error rate. This feature can be utilized by dentists to better track patient dental behavior.
@@ 8.6.3. Tracking other activities
In our study, we were only able to track 27 activities. However, our system has the potential to recognize other activities that involve body pose/movements on the upper body and face. Researchers can potentially replicate our system and customize the frameworks to detect the activity of their interest. For instance, this system can be easily used to automatically track and log the duration and types of the user’s exercise routines.
@@ 8.7. Limitations and Future Work
8.7.1. Scope of Activities in the Dataset
ActSonic recognizes 27 distinct everyday activities within its dataset. However, certain activities in the dataset exhibit variability in execution. For example, actions like yawning or pouring can vary based on contextual factors, affecting system performance in real-world settings. Addressing this diversity might benefit from a larger dataset and a foundational deep learning feature extractor.
@@ 8.7.2. Usage of Differential Echo Profile Only
8.7.3. Lack of Multi-Label/Concurrent Activity Detection
Real-world scenarios involve concurrent activities, a challenge yet to be explored in wearable technology. We aim to explore multi-label classifiers leveraging our system’s groundbreaking performance.
8.7.4. Lack of External Contextual Detection
Our focus on upper body motion doesn’t consider the contextual environment. Integrating GPS or camera data could aid in understanding environmental affordances, and improving activity detection.
8.7.5. Lack of Access to Fine-Grained Hand or Head Movements
Our system currently captures upper-body motions only. Incorporating IMU sensors for head movements or additional acoustic sensors for precise hand tracking could enhance our approach through a multi-modal setup.

### v2

@@ 8.7. Potential Real-world Application
The promising performance of ActSonic recognizing 27 activities in the wild using low-power and minimally obtrusive glasses will significantly lower the barriers to logging everyday activities. It would further create opportunities for many downstream applications that are based on tracking one or multiple types of activities. Here we list a few sample applications :
@@ 8.7.2. High-Resolution Behavior Data in the Wild for Health
ActSonic can track 27 everyday activities, many of which are related to health behaviors. Automatically logging these health-related activities can potentially provide researchers and clinical physicians with an opportunity to better understand a user’s activities in the wild for health purposes. For instance, eating and drinking behavior can potentially be used to analyze the user’s eating and drinking routines related to eating disorders and body hydration levels. Brushing teeth, flossing, and rinsing the mouth can be recognized with a very low error rate. This feature can be utilized by dentists to better track patient dental behavior.
@@ 8.7.3. Tracking Other Activities
In our study, we were able to track 27 activities. However, our system has the potential to recognize other activities that involve movements on the upper body and face. Researchers can potentially replicate our system and customize the frameworks to detect the activity of their interest. For instance, this system can be easily used to automatically track and log the duration and types of the user’s exercise routines.
@@ 8.8. Limitations and Future Work
8.8.1. Scope of activity information in the dataset
ActSonic recognizes 27 distinct everyday activities within its dataset. However, certain activities in the dataset exhibit variability in execution. For example, actions like yawning or pouring can vary based on contextual factors, which can affect system performance in real-world settings. Addressing this diversity might benefit from a larger dataset and a foundational deep learning feature extractor. Moreover, the activity recognition pipeline of ActSonic lacks external contextual information. Integrating GPS or motion data for fine-grained head or hand movements could aid in understanding environmental affordances and improving activity detection.
@@ 8.8.2. Usage of differential echo profile only
8.8.3. Multi-Label or concurrent activity detection
Real-world scenarios involve concurrent activities, a challenge yet to be explored in wearable technology. We aim to explore multi-label classifiers leveraging our system’s superior performance.